OpenStudy (anonymous):

The sum of an arithmetic series is 1356 and the first term is -1. If the common difference is 5, how many terms are in the sequence?

OpenStudy (ash2326):

@katiebugg Sum of n terms is given 1356 First term \(a_1=-1\) and common difference=5 we know that sum to n terms is given as \[S_n=\frac{n}{2}(a_1+a_n)\] \(a_n\) is the last term Do you get till this?

OpenStudy (anonymous):

yess

OpenStudy (ash2326):

we know \(a_n\) is given as \[a_n=a_1+(n-1)d\] d=common difference n=no. of terms so sum will become \[S_n=\frac n2 (a_1+a_1+(n-1)d)\] now pluigin the numbers here and post what do you get, will you?